Georgia is jogging with a velocity of 6 m/s when she accelerates at 2 m/s/s for 5 seconds. How fast is Georgia running now?
1
Expert's answer
2020-11-10T06:59:04-0500

The velocity can be found by


"v_f=v_i+at=6+2\cdot5=16\text{ m/s}."